OpenStudy (anonymous):

Need this last one and I'm set. A washer and a dryer cost $1000 combined. The cost of the washer is three times the cost of the dryer. What is the cost of the dryer?

OpenStudy (mathstudent55):

The cost of the dryer = d The cost of the washer = w Can you express the cost of the washer in terms of the dryer? w = ?

OpenStudy (anonymous):

you set it up like this: Washer=W, Dryer=D 1000=W+D W=3D so 1000=D+3D or 1000=4D divide both sides by 4 and we get D=250 so Dryer = $250 and Washer = $750

OpenStudy (anonymous):

Thank you chanellos. I'll keep that in mind :)